OpenStudy (lrki):

Please Help... How can I simplify this expression 2 √-24+6 using the imaginary unit i

OpenStudy (anonymous):

Alright lets work on this shall we? @Lrki

OpenStudy (anonymous):

So we have \[2\sqrt{-24+6}\] right?

OpenStudy (lrki):

Yes

OpenStudy (anonymous):

or is it \[2\sqrt{-24}+6\]

OpenStudy (anonymous):

The 1st one or the second one?

OpenStudy (lrki):

The 1st

OpenStudy (anonymous):

Okay so now we have \[2\sqrt{24+6}\]

OpenStudy (anonymous):

Sorry with a negative in front.

OpenStudy (lrki):

My bad it's the second one.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Now what we need to do is make the 24 into a positive and multiply it with -1 like this: \[2\sqrt{24(-1)}+6\]

OpenStudy (lrki):

Ok

OpenStudy (anonymous):

Do you see what I did?

OpenStudy (lrki):

Yes

OpenStudy (lrki):

Would the answer be 6+4i √6

OpenStudy (anonymous):

Now you want to separate the -1 into a separate square root \[2\sqrt{24}\sqrt{-1}+6\]

OpenStudy (anonymous):

Well lets work it out I don't know the answer yet.

OpenStudy (anonymous):

Now \[\sqrt{-1}\]

OpenStudy (anonymous):

is equal to \[i\]

OpenStudy (lrki):

I got the answer thanks anyways.

OpenStudy (anonymous):

Oh haha okay xDD

OpenStudy (anonymous):

Sorry for the time.

OpenStudy (anonymous):

Hope I helped anyway.

OpenStudy (anonymous):

Have a great day!